OpenStudy (anonymous):

Can somebody explain how my solutions manual went from (- 2 + √-20)/6 to - 2/6 + (2i√5)/6? I don't see how they came up with the 2i*5^2=-20^2. i = -1^2. Equation editor is not working for me..

OpenStudy (anonymous):

The square root of -1 is i. sqrt(-20) = sqrt(-1 * 4 * 5) Take out the -1 and the 4 and you get 2i * sqrt(5)

OpenStudy (anonymous):

\[\sqrt{-20} = \sqrt{-1\times 4\times 5} \]\[= \sqrt{-1}\times \sqrt{4} \times \sqrt{5}\]\[=i \times 2 \times \sqrt{5}\]\[=2i\sqrt{5}\]

OpenStudy (anonymous):

2i*sqrt(5) would equal -10, would'nt it?

OpenStudy (anonymous):

No. It equals \[2i \sqrt{5}\]

OpenStudy (anonymous):

I'm, confused...

OpenStudy (anonymous):

When you have the square root of a negative number, you need to break it up like polpak did. You can factor out the square root of -1 which is always equal to i.

OpenStudy (anonymous):

It would also equal \(2\sqrt{-5}\) or \(i\sqrt{20}\)

OpenStudy (anonymous):

But the simplest form is the one I gave you originally.